OpenStudy (maggy14145):

What is the average rate of change for the relation shown below between x = -2 and x = 2?

OpenStudy (maggy14145):

http://cds.flipswitch.com/tools/asset/media/385879

OpenStudy (maggy14145):

PLEASE HELP ME @ShadowLegendX

OpenStudy (shadowlegendx):

Do you know how to make a table for a relation?

OpenStudy (maggy14145):

no not really

jimthompson5910 (jim_thompson5910):

draw a vertical line through -2 on the x axis. Where does this vertical line cross the parabola?

OpenStudy (maggy14145):

-2,8

jimthompson5910 (jim_thompson5910):

use parenthesis (-2,8)

jimthompson5910 (jim_thompson5910):

do the same for x = 2

OpenStudy (maggy14145):

(2,0)

jimthompson5910 (jim_thompson5910):

so they want you to find the slope of the line through the two points (-2,8) and (2,0) Slope Formula \[\Large m = \frac{y_2 - y_1}{x_2 - x_1}\]

OpenStudy (maggy14145):

-2 is the answer

OpenStudy (maggy14145):

but it says to put the answer in fraction form..

jimthompson5910 (jim_thompson5910):

yes -2 is the answer. If your teacher wants it in fraction form, then write it as -2/1

jimthompson5910 (jim_thompson5910):

since x/1 = x

OpenStudy (maggy14145):

ok and this is another question. the one about the banquet that I got mixed up in

OpenStudy (maggy14145):

A banquet hall charges $9.50 per person if the group size is 25 people or less and a $650 flat rate if the group size is more than 25 people. Which function represents the cost of hosting a party at that banquet hall?

OpenStudy (maggy14145):

f(x)=9.50x+650

OpenStudy (maggy14145):

f(x)=650-9.50x^2

OpenStudy (maggy14145):

f(x)=650(x-25)+9.50

jimthompson5910 (jim_thompson5910):

the attachment isn't working. I'm not sure why

OpenStudy (maggy14145):

hold on I'm trying to find a way to get it on here

OpenStudy (maggy14145):

ok I can't get the picture but this is what it looks like

OpenStudy (maggy14145):

f(x)={ 9.50x 0 <= x <= 25 650 25

jimthompson5910 (jim_thompson5910):

Sounds like it looks like this \[\Large f(x) = \begin{cases}9.50x \ \ \ \text{ if } 0 \le x \le 25\\650 \ \ \ \text{ if } x > 25\end{cases}\] is my assumption correct?

OpenStudy (maggy14145):

yes pretty much

jimthompson5910 (jim_thompson5910):

so the first part says "if x is between 0 and 25, then the function is f(x) = 9.50x" the second part says "if x is larger than 25, then f(x) = 650"

jimthompson5910 (jim_thompson5910):

f(x) changes depending on what x is

OpenStudy (maggy14145):

okay

jimthompson5910 (jim_thompson5910):

does that make sense?

OpenStudy (maggy14145):

yes so far

jimthompson5910 (jim_thompson5910):

Do you see how that models the cost function? Do you see how it takes care of both cases?

OpenStudy (maggy14145):

yes I can see that

jimthompson5910 (jim_thompson5910):

so that would be the answer

OpenStudy (maggy14145):

so it would be the second one?

jimthompson5910 (jim_thompson5910):

The whole thing \[\Large f(x) = \begin{cases}9.50x \ \ \ \text{ if } 0 \le x \le 25\\650 \ \ \ \text{ if } x > 25\end{cases}\] is considered one function. It's called a piecewise function.

jimthompson5910 (jim_thompson5910):

it's comprised of two different pieces to make it one big function

jimthompson5910 (jim_thompson5910):

Think of taking apart 2 cars and joining them together to form one bigger car

OpenStudy (maggy14145):

ohhh ok I am aware of piecewise functions. this makes a lot more sense now

OpenStudy (maggy14145):

you must love working on math since you've been helping me for the past few hours haha

jimthompson5910 (jim_thompson5910):

it's a lot of fun lol

OpenStudy (maggy14145):

I'm trying to finish an assignment but I'm just so confused on a ton of them I don't even know where to begin :(

jimthompson5910 (jim_thompson5910):

ok what other question did you have?

OpenStudy (maggy14145):

The parabola graphed below represents a side view of a parabolic microphone dish. The directrix is drawn in as well. The end of the microphone is placed on a feedhorn that will need to be placed at the focus of the parabola to pick up the sound. The scale of the graph is in inches. Based on the graph, the engineers should make the feedhorn _____ inches long. Enter your answer as the number that correctly fills in the blank in the previous sentence.

OpenStudy (maggy14145):

http://cds.flipswitch.com/tools/asset/media/385896

OpenStudy (maggy14145):

would the correct answer be 4?

jimthompson5910 (jim_thompson5910):

To be honest, I'm not exactly sure what they're asking. It sounds like they want you to place the feedhorn at the focus so if you want it to connect to the parabola, then yes, it would be 4 inches. This is because the distance from vertex to directrix is also 4 inches.

OpenStudy (maggy14145):

ok its fine. next question.Given f(x) = 2x2 + 2x and g(x) = x + 1, find the function (f/g)(x).

OpenStudy (maggy14145):

I know the answer is 2x but does x= 0 or -1 ??

jimthompson5910 (jim_thompson5910):

the denominator is x+1 it cannot be zero. If it were, then x+1 = 0 leads to x = -1 If x = -1,then the denominator is 0. To avoid dividing by zero, we must add on the restriction that \(\Large x \ne -1\)

OpenStudy (maggy14145):

Which exponential model described below has the smallest y-intercept?

OpenStudy (maggy14145):

y=6^x+4

OpenStudy (maggy14145):

a model describes a sample that doubles every hour for x hours.

OpenStudy (maggy14145):

http://cds.flipswitch.com/tools/asset/media/385816

OpenStudy (maggy14145):

http://cds.flipswitch.com/tools/asset/media/385817

jimthompson5910 (jim_thompson5910):

when you write `y=6^x+4` do you mean \(\Large 6^{x}+4\) or do you mean \(\Large 6^{x+4}\)

OpenStudy (maggy14145):

the first one u put

OpenStudy (maggy14145):

x is the exponent of 6

jimthompson5910 (jim_thompson5910):

what are you thinking on this one? were you able to eliminate a few options?

OpenStudy (maggy14145):

I removed the t-chart

jimthompson5910 (jim_thompson5910):

why did you remove the t-chart ? what's your thinking on that decision?

OpenStudy (maggy14145):

wait hold on I think that actually might be the correct answer

OpenStudy (maggy14145):

since it has the smallest x-value which is 0

jimthompson5910 (jim_thompson5910):

the y intercept happens when x = 0. They want you to find the smallest y value when x = 0. So you'll compare the different y intercepts of the functions

OpenStudy (maggy14145):

so the y-int would be 4

jimthompson5910 (jim_thompson5910):

on the t-chart? or are you looking at the graph now?

OpenStudy (maggy14145):

on the t chart

jimthompson5910 (jim_thompson5910):

look at where x = 0. What is the paired y value here?

OpenStudy (maggy14145):

64

jimthompson5910 (jim_thompson5910):

so the y intercept here is 64 ie, the point (0,64)

OpenStudy (maggy14145):

correct

jimthompson5910 (jim_thompson5910):

what's the y intercept of the graph?

OpenStudy (maggy14145):

(0,3)

OpenStudy (maggy14145):

oh this looks like the correct answer

OpenStudy (maggy14145):

yes! that was the correct answer

jimthompson5910 (jim_thompson5910):

yes it is the correct answer. If you plugged in x = 0 into 6^x + 4, you'd find that y = 5

OpenStudy (maggy14145):

2(10^3x ) = 24 how can I make this into a logarithm

jimthompson5910 (jim_thompson5910):

The equation given to you is this? or no? \[\Large 2\left(10^{3x}\right) = 24\]

OpenStudy (maggy14145):

yes

jimthompson5910 (jim_thompson5910):

the first thing I would do is divide both sides by 2

jimthompson5910 (jim_thompson5910):

what do you get when that happens?

OpenStudy (maggy14145):

(10^3x)=12

jimthompson5910 (jim_thompson5910):

yes. Use parenthesis to show that 3x is in the exponent 10^(3x) = 12

jimthompson5910 (jim_thompson5910):

what comes next?

OpenStudy (maggy14145):

Im confused. would it be log12=

jimthompson5910 (jim_thompson5910):

that's part of it

jimthompson5910 (jim_thompson5910):

you apply logs to both sides or convert to log form

OpenStudy (maggy14145):

so is log12 correct?

OpenStudy (maggy14145):

as part of the answer?

jimthompson5910 (jim_thompson5910):

\[\Large 10^{3x} = 12\] turns into \[\Large 3x = \log(12)\] where the log is base 10

OpenStudy (maggy14145):

oh ok so u just flipped the log to the other side?

OpenStudy (maggy14145):

is that the most simplified version of the answer?

jimthompson5910 (jim_thompson5910):

does your teacher want you to solve for x?

OpenStudy (maggy14145):

they just want the most simplest form of that equation

OpenStudy (maggy14145):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

so it sounds like they want to solve for x. So you can divide both sides by 3 or multiply both sides by 1/3

OpenStudy (maggy14145):

x =log(4)